# Are Bus-Mounted Edge Servers Feasible?

### v3

@@ III-A Datasets
The datasets span over a wide range of the Shanghai city, but we select the center of the map and partition it into grid cells with a side length of $200$ meters. The selected region is a square area that covers $72\times 72$ grids, containing $937$ base stations, $2,100$ buses, and $26$ million taxi GPS records. The high density of the base stations, buses, and taxis in this area allows us to build a thorough understanding of the coverage of buses and base stations from the trace data. We divide the records in the trace into discrete time slots (e.g., 30 seconds each). For each time slot, we retrieve the locations of the base staions, buses, and taxis, and evaluate how many grids or taxis (demand points) can be covered by the base stations or buses, given a service radius $R$ of these potential servers.
@@ III-B1 Coverage of Base Stations
The coverage of the base stations on the demand points on April 1st in $2018$ is shown in Fig. 2. We select different portions of base stations randomly to deploy edge servers, and also vary their service radius. The varying percentages of selected base stations are $10\%,20\%,30\%,40\%,50\%,60\%]$, and the varying radii are $[0.4,0.6,0.8,1.0,1.2]$ km.
@@
Fig. 2(a) illustrates the area coverage of the base stations, showing that the existing base station layout can cover most of the area. For instance, with a service radius of $1$ km, $20\%$ of base stations yields nearly $90\%$ area coverage. Fig. 2(b) shows the point coverage, and the result also indicates a near-perfect coverage when the server computation capacity limit is ignored. However, Fig. 2(c) shows much worse results when we adopt a more realistic setting by limiting the server capacity to $50$ requests per time slot. The coverage decreases remarkably for all radii, which indicates that the capacity is in fact the primary bottleneck rather than the radius. It is worth noting that deploying edge servers with sufficient capacity to cover all demands all the time is irrational, as such capacity during normal periods may be wasted.
@@ III-B2 Coverage of Buses
The results show that buses also demonstrate desirable performance in terms of the three coverage metrics. The area and point coverage of buses, although smaller than base stations, can also reach over $50\%$ for certain server percentages and radii as shown in Fig. 3(a) and Fig. 3(b). Under more realistic scenarios shown in Fig. 3(c), the value of buses becomes pronounced. With a service radius greater than $0.8$ km and $50\%$ (approx. $1,070$) random buses equipped with edge servers, over $40\%$ of the demand points can be covered. Note that such point coverage ratio is based on the corresponding $27\%$-area coverage ratio. This indicates that the buses always go through high-traffic zones, which have a great potential in acting as edge servers and compensating for the server capacity deficit of base station-based edge servers.
@@ V-B Performance Results
Fig. 6 shows a comparison of the demand volume covered by base stations and buses during the evening peak hour from 18:00 to 20:00, to show the effect of deploying both base station and bus-based edge servers. Fig. 6(a) shows the demand volume of the end users without any edge servers. Fig. 6(b) shows the remaining demand after deploying $300$ random base stations, and Fig. 6(c) shows the result after introducing additional $200$ heuristically selected buses on that basis. Both the base stations and buses have a service radius of $0.8$ km and a capacity of $50$. The results reveal that while base stations handle most of the demands, some downtown areas still exhibit high load (the grids with red color), yet the buses effectively eliminate these remaining demand hotspots. This indicates that bus-mounted edge servers can cooperate with the base stations well, effectively handle the bursty traffic overwhelming the computation capacity of base stations during rush hours, and thereby improve the system computation elasticity and overall quality of service.

### v4

@@ III-A Datasets
The datasets span over a wide range of the Shanghai city, but we select the center of the map and partition it into grid cells with a side length of $200$ meters. The selected region is a square area that covers $72\times 72$ grids, containing $934$ base stations, $2,143$ buses, and over $26$ million taxi GPS records. The high density of the base stations, buses, and taxis in this area allows us to build a thorough understanding of the coverage of buses and base stations from the trace data. We divide the records in the trace into discrete time slots (e.g., 30 seconds each). For each time slot, we retrieve the locations of the base staions, buses, and taxis, and evaluate how many grids or taxis (demand points) can be covered by the base stations or buses, given a service radius $R$ of these potential servers.
@@ III-B1 Coverage of Base Stations
The coverage of the base stations on the demand points on April 1st in $2018$ is shown in Fig. 2. We select different portions of base stations randomly to deploy edge servers, and also vary their service radius. The number of randomly selected base stations is varying from $100$ to $900$, and the radius is varying from $0.4$ km to $1.2$ km.
@@
Fig. 2(a) illustrates the area coverage of the base stations, showing that the existing base station layout can cover most of the area. For instance, with a service radius of $0.8$ km, $500$ base stations yield over $90\%$ area coverage. Fig. 2(b) shows the point coverage, and the result also indicates a near-perfect coverage when the server computation capacity limit is ignored. However, Fig. 2(c) shows much worse results when we adopt a more realistic setting by limiting the server capacity to $50$ requests per time slot. The coverage decreases remarkably for all radii, which indicates that the capacity is in fact the primary bottleneck rather than the radius. It is worth noting that deploying edge servers with sufficient capacity to cover all demands all the time is irrational, as such capacity during normal periods may be wasted.
@@ III-B2 Coverage of Buses
The results show that buses also demonstrate desirable performance in terms of the three coverage metrics. The area and point coverage of buses, although smaller than base stations, can also reach over $50\%$ for certain server percentages and radii as shown in Fig. 3(a) and Fig. 3(b). Even under more realistic scenarios shown in Fig. 3(c), the buses still cover a considerable percentage of the user demands. With a service radius of $0.8$ km and $900$ random buses equipped with edge servers, nearly $40\%$ of the demand points can be covered. This indicates that the buses always go through high-traffic zones, which have a great potential in acting as edge servers and compensating for the server capacity deficit of base station-based edge servers.
@@ V-B Performance Results
Fig. 6 shows a comparison of the demand volume covered by base stations and buses during the evening peak hour from 18:00 to 20:00, to show the effect of deploying both base station and bus-based edge servers. Fig. 6(a) shows the demands of the end users without any edge servers. Fig. 6(b) shows the remaining demands after deploying $300$ random base stations, and Fig. 6(c) shows the result after introducing additional $200$ heuristically selected buses on that basis. Both the base stations and buses have a service radius of $0.8$ km and a capacity of $50$. The results reveal that while base stations handle most of the demands, some downtown areas still exhibit high load (the grids with red color), yet the buses effectively eliminate these remaining demand hotspots. This indicates that bus-mounted edge servers can cooperate with the base stations well, effectively handle the bursty traffic overwhelming the computation capacity of base stations during rush hours, and thereby improve the system computation elasticity and overall quality of service.